### v2

@@ Abstract
We begin with a retrospective analysis of three major U.S. airline mergers and document the sensitivity of the findings, particularly questioning whether market conditions evolve similarly for treated and control markets. We then develop a structural model that clarifies this and other assumptions implicit in retrospective analyses and separates efficiency gains from increases in firms’ conduct. Using only pre-merger data, we propose a reduced-form approach that leverages exogenous changes in market structure to forecast merger effects. Finally, we use structural prospective merger simulations with our other estimates for a comprehensive evaluation. This bridging of approaches uncovers a fundamental tension: either efficiency gains were limited or, if they were significant, they were accompanied and offset by coordinated effects.
@@ 1 Introduction
Merger evaluation remains one of the most challenging problems in antitrust policy because of the difficulties separating efficiency gains from increased market power. This literature has developed two distinct methodologies: retrospective analyses that examine post-merger outcomes by applying event study methods like difference-in-differences (DiD) to pre- and post-merger data (e.g., Kim and Singal, 1993; Focarelli and Panetta, 2003; Hastings, 2004; Ashenfelter and Hosken, 2010; Ashenfelter et al., 2015; Carlton et al., 2019), and prospective analyses that predict merger effects using only pre-merger data (e.g., Nevo, 2000; Nevo and Whinston, 2010). There are a few important exceptions to this classification (Baker, 1999; Pesendorfer, 2003; Davis, 2005; Miller and Weinberg, 2017; Igami and Sugaya, 2022; Bruegge et al., 2024), which underpin our work. For a more comprehensive overview of mergers see Whinston (2006); Kwoka, Jr. (2014); Asker and Nocke (2021) and Kaplow (2024).
@@
We aim to fill this gap by developing a unified approach that bridges these methodologies, demonstrating its value through an analysis of three major airline mergers: Delta-Northwest (2008), United-Continental (2010), and American-US Airways (2013). Our empirical framework operates along two distinct methodological dimensions, assuming that mergers and market structures are exogenous. The first dimension distinguishes between analyses using pre- and post-merger data versus those using only pre-merger data. The second separates reduced-form from the structural approach, creating four complementary analyses (see the four quadrants in Table 1) that provide a comprehensive understanding of merger effects. In the process, we show how insights from one approach can inform and strengthen the other.
@@
Reduced Form Retrospective Analysis Market Structure Analysis
@@
Our paper follows a systematic analysis that sequentially integrates reduced-form and structural approaches. We begin with a reduced-form retrospective analysis of pre- and post-merger data (Section 3), then evaluate the robustness of this approach concerning the time trends to capture possibly declining demand. Next, we develop an equilibrium model with linear market-level demand that links pricing behavior to mergers (Section 4). This model allows us to interpret the DiD estimation equation from Section 3 as a reduced-form expression of our structural pricing equation, explicitly delineating the economic assumptions that underpin the DiD strategy. Then, we develop a regression-based approach that relies on the exogenous variation in market structure in only the pre-merger data to forecast merger effects (Section 5). We conclude by implementing a structural merger simulation that employs only the pre-merger data and uses estimated effects of mergers on prices from our Sections 3 and 5 to determine the post-merger marginal costs and use it to evaluate the reduced form estimates under different post-merger strategic conduct based on the findings in Section 4 (Section 6).
The US airline industry provides an ideal setting for our analysis. It offers rich, comprehensive data spanning ticket prices, passenger numbers, seat capacity, and flight schedules from multiple authoritative sources: the Department of Transportation’s Origin and Destination Survey for ticket prices and passenger numbers, the Bureau of Transportation Statistics’ T100 database for seat capacity, and the Official Airline Guide (OAG) data for flight schedules. Moreover, the industry’s extensive scrutiny in both academic research (e.g., Borenstein, 1990; Kim and Singal, 1993; Olley and Town, 2018; Carlton et al., 2019; Das, 2019; Li et al., 2022; Eizenberg and Zvuluni, 2024; Olsen et al., 2024; Remer and Orchinik, 2024) and legal proceedings (e.g., U.S. District Court, MA, 2023) provides crucial institutional context and validation points for our methodology
@@
Next, to explore these findings and make explicit the economic assumptions underlying the DiD framework, we develop a structural framework in Section 4 such that the implied reduced form price equation corresponds to the one used in retrospective analysis. DiD is a program evaluation that requires a set of assumptions about airline pricing to determine the effect of a merger, and our structural model clarifies those hypotheses. To this end, we build on Bresnahan (1982) and propose a structural model that allows merger-specific coordinated effects (Porter, 2020), treating air travel as a homogeneous product with market-level quantity-weighted average prices.
@@
Our equilibrium model uncovers three critical insights in the implicit assumptions of standard retrospective analyses. First, consistent DiD estimation requires that mergers not alter firm conduct and that time-varying unobservable “demand shocks” evolve similarly across treated and control markets. Second, the error term in the DiD specification contains unobservable demand and supply shocks. While the cost shocks may reasonably be uncorrelated with merger decisions, demand shocks could systematically differ between treated and control markets—for instance, if firms are more likely to merge in markets with declining demand—and biasing the DiD estimates. Third, our estimates suggest that mergers affect firm conduct differently across market structures, indicating both unilateral and coordinated effects that the retrospective analysis might miss.
@@
Next, we consider the scenario typically faced by the antitrust authorities and the courts: evaluating mergers using only pre-merger data. In Section 5, we develop a regression-based approach that leverages exogenous variations in market structure pre-merger to predict post-merger effects using only pre-merger data. Building on methodology pioneered by Baker (1999), this approach analyzes how prices vary with different market structures in the pre-merger period to forecast merger-induced changes. The variation in market structure offers insights into how similar structural changes might affect post-merger outcomes. We find that transitions from duopoly to monopoly are associated with modest but consistent fare increases (ranging from 2.9% to 6.2%) and substantial capacity reductions (15.7% to 27.6%). Conversely, markets transitioning from monopoly to duopoly show symmetric effects: fare decreases of 3.0% to 7.7%, and large capacity increases of 31.5% to 38.5%. While intuitive and simple, this model also captures efficiency induced by past changes in market structure. When a market evolved from a triopoly to a duopoly, the remaining airlines likely optimized their operations in response—and those efficiencies are present in the observed price changes.
@@
Finally, in Section 6, we implement an equilibrium-based merger simulation (Nevo, 2000) that considers different scenarios for post-merger cost efficiencies and conduct. In particular, we consider a random utility discrete choice model of demand with differentiated products and Bertrand competition where firms choose prices. Following Verboven (1996), we employ a nested logit demand model to distinguish between nonstop and connecting services, allowing for rich substitution patterns across products and carriers. Our merger simulation analysis reveals that predicted price effects depend critically on assumptions about post-merger cost efficiencies and firms’ conduct. In particular, we explore how post-merger prices depend on efficiency gains and changes in competitive conduct. Under standard competition assumptions and maximum efficiency gains (where merged entities adopt the lower costs of merging firms), the model predicts price decreases ranging from 2.9% to 5.1%. However, this finding presents a puzzle when compared to our earlier findings of stable post-merger prices: either efficiency gains were limited or significant efficiencies were offset by an increase in coordinated effects.
Thus, our bridging framework reveals a crucial tension in evaluating airline mergers that traditional approaches miss. As post-merger prices remain stable, there are two possible interpretations: either the merger generated substantial efficiency gains but also increased market power significantly, or both efficiency gains and competitive effects were modest. This tradeoff creates a strategic dilemma for merging airlines - they can either claim large efficiency gains but must also accept the evidence of increased market power or argue for modest competitive effects. However, then it weakens the case for the merger based on efficiency. By iterating between reduced-form and structural approaches, our framework makes this tradeoff explicit, providing authorities with a more complete toolkit for merger evaluation. Moreover, our finding that mergers primarily affect capacity rather than prices suggests that antitrust authorities should broaden their focus beyond price effects alone.
@@
Our paper’s contributions to the literature are both methodological and empirical. Building on the extensive merger analysis literature cited throughout this introduction and the rest of the paper, we integrate existing reduced-form retrospective and structural prospective analyses, drawing insights from each while addressing their limitations. Specifically, we demonstrate how retrospective studies that use post-merger data to measure price effects (Ashenfelter and Hosken, 2010) and prospective approaches that use pre-merger data and focus on structural estimation of efficiency gains (Farrell and Shapiro, 1990) can be integrated to provide more robust merger analysis, under the assumption of exogenous mergers. [^2] This assumption is nearly universal in the empirical merger literature, underlying both retrospective analyses and structural merger simulations, with notable exceptions being Dafny (2009); Aguirregabiria and Ho (2012); Ciliberto et al. (2021); Li et al. (2022), and Bontemps et al. (2023) among others.
@@ 2 Data and Descriptive Analysis
We focus on three major mergers in the airline industry. Delta and Northwest Airlines announced their merger plan on April 14, 2008; the shareholders approved it on September 26, 2008, and the US Department of Justice (DOJ) approved it on October 29, 2008. United and Continental announced their merger plan on May 2, 2010, gaining shareholder approval on September 17, 2010. The merger was completed on October 1, 2010, forming United Continental Holdings, although the full operational integration was completed only in mid-2012. The merger between American Airlines and US Airways was announced on February 14, 2013. This merger faced initial opposition from the DOJ, which filed a lawsuit on August 13, 2013, to block the proposed merger. However, a settlement was reached on November 12, 2013, allowing the merger to be completed on December 9, 2013.333While (Kim and Singal, 1993) examine price effects at both merger announcement and implementation (approval) dates, our analysis focuses exclusively on the pre- and post-approval periods. This approach allows us to focus on the real market effects following regulatory clearance rather than the anticipatory effects, providing a more focused examination of realized competitive outcomes.
@@
Next, we discuss the selection of “treated” and “control” markets, an approach the US court system accepts when evaluating merger cases. [^4] In our analysis, we also consider the sample of markets connecting the top 50 Metropolitan Statistical Areas (MSAs), a standard approach widely used in the literature. For example, Berry and Jia (2010) examine airports located in medium to large metropolitan areas with at least 850,000 people in the year 2006, which roughly corresponds to the top 50 MSAs. For each merger, we classify a city pair market as a “treated” market if both merging parties provided nonstop service before the merger. [^5] A carrier is considered ‘nonstop’ in a city route if it conducts 10 nonstop operations (five round-trips). Furthermore, we separately focus on 2-to-1 markets, the duopoly markets, where only the merging parties provide nonstop service, and 3-to-2 markets, the triopoly markets, where one other carrier provides nonstop service besides the two merging parties.
@@
Likewise, we classify a market as a “control” market if the merger does not affect its market structure. In other words, in such a market, only at most one carrier can be the merging party, so if it is a duopoly or a triopoly pre-merger, it remains the same post-merger. Thus, the control markets have the same number of nonstop carriers as the overlap routes in the pre-merger period. As with the treated market, we also restrict our attention to control markets where, whenever applicable, the merging party does not have a substantial connecting presence.
@@ 3 Retrospective Analysis
We use the event study approach to separately assess the impact of three legacy airline mergers (Delta-Northwest, United-Continental, and American-US Airways) on three key economic variables: average fares, passenger volume, and seat availability.
The analysis compares the mean of each variable in control markets before and after a merger and compares that difference with the difference in means in treated markets. In particular, we estimate the following regression specification to determine the average treatment on the treated under the DiD assumptions:
@@ 3.1 Estimation Results
Our parameter of interest is $\beta^{\texttt{DiD}}$, which, under the parallel trends assumptions, is the effect of a merger on the outcome variables in the treated markets. The estimation results from (1) are shown in Table 3. Panel A presents estimates using the fixed effects method (mean deviation). We estimate that the average prices decreased by 4.02% (DL-NW), 3.54% (UA-CO), and 7.50% (AA-US). The passenger volume increased by 2.84% (DL-NW), 5.87% (UA-CO), and 9.19% (AA-US). [^6] This estimate is a weighted average of the parameter estimates across markets and time (de Chaisemartin and D’Haultfœuille, 2020) and should be interpreted as a percentage change in the outcome variable. The percentage changes for $\beta^{\texttt{DiD}}$ is given by $(\exp(\beta^{\texttt{DiD}})-1)\times 100$ (Halvorsen and Palmquist, 1980). Almost all of these estimates are statistically significant at 1%-level, except that the estimated effect on passenger volume following the DL-NW merger is noisy. Most notably, the number of offered seats, i.e., the allocated capacities, expanded by 24.11% (DL-NW), 29.44% (UA-CO), and 16.53% (AA-US) - all statistically significant.
@@ 3.2 Robustness Analysis
The most natural and simple way to check the robustness is to run first differences instead of fixed effects, as done in Kim and Singal (1993). The first difference and fixed effect methods should yield similar estimates. [^8] They yield the same estimates if there is only one period before and one period after the mergers. However, they differ in their identifying assumptions, which in the presence of time trends may result in different estimates (Wooldridge, 2010). In particular, the fixed effects method requires strict exogeneity, i.e., $\text{Treated}_{mt}\times\text{Post-Merger}_{mt}$ is uncorrelated with $\varepsilon_{mt}$ across all periods. In contrast, first differences only require that the variable is uncorrelated with the change in $\varepsilon_{mt}$ immediately before and after the merger. The variable $\text{Treated}_{mt}\times\text{Post-Merger}_{mt},$ changes value only in a single quarter, suggesting that the first differences assumption is less demanding than the fixed effect method. Furthermore, in a DiD framework, the first differences have an additional advantage: the constant term captures any differential linear trends between treatment and control markets, slightly relaxing the parallel trends assumption.
The results from using the first differences to estimate (1) are shown in Table 3 Panel B. The contrast with our fixed effects estimates is striking in three key dimensions.
First, while fixed effects estimation suggests that mergers lead to a 5.6% decrease in prices, the first differences approach suggests prices increase slightly. However, they are measured with noise, except for the UA-CO merger, for which the price increases by 4.5%. Second, the effects of the merger on passenger volume are significantly different. The DL-NW merger shows a substantial 21.8% decrease in passengers using first differences, compared to a small positive effect using the fixed effects method. Only AA-US maintains a positive passenger effect (27.5% increase) in first differences. Third, while capacity effects remain positive and significant across both specifications, their magnitude is notably smaller under the first differences approach. For example, the UA-CO capacity effect drops from 25.8% to 3.7%.
The first differences method implicitly allows for different time trends between treatment and control groups through the constant term. In the airline context, such differential trends could reflect market-specific changes in local economic conditions or evolving travel patterns. The differences in the estimates using the two methods highlight the importance of capturing time trends for merger analysis.
The stark differences between fixed effects and first differences estimates raise concerns about whether the parallel trends assumption—fundamental to DiD analysis—holds in our setting. To investigate this issue more rigorously, we conduct pre-trend analyses using the event study approach.
@@
Our pre-trend analysis in Figure 2 indicates that the parallel trends assumption critical to DiD methodology is violated in this context. One approach to address this violation is including market-specific time trends in our model. In airline markets, such trends reflect gradual changes in local economic conditions or evolving travel patterns, specifically on the demand side. However, incorporating these trends presents methodological challenges, particularly in choosing the appropriate functional form.
@@
Table 4 presents these results. Notably, incorporating time trends substantially alters the main findings Table 3 Panel A: price decreases, and passenger volume increases either reverse or become statistically insignificant. This pattern holds across different trend specifications, suggesting the initial findings may have been driven by pre-existing market trajectories rather than merger effects.
@@
While our linear trend specification imposes strong assumptions about market evolution, specifically that trends are linear and homogeneous within groups, we validate our findings using more flexible methods. In Appendix B.6, we demonstrate that the synthetic DiD approach, which better captures the complex dynamics of airline markets, produces qualitatively similar results. This consistency across methodologies strengthens our confidence in the robustness of our findings. [^10] Specifically, The estimates from synthetic DiD are in Appendix B.6 Table B.5 Panel B and are similar to the first-difference results reported in Table 3 Panel B for prices and passenger volume. Our findings are similar to Remer and Orchinik (2024), who uses the synthetic DiD method and find that mergers lead to higher prices.

### v3

@@ Abstract
We propose an ensemble approach to evaluate mergers that combines retrospective and prospective modeling for a more reliable analysis. We begin with a retrospective analysis based on an event study of three major U.S. airline mergers and document the fragility of the findings. We then develop a structural model that nests the event study in its reduced form, clarifying the implicit assumptions in retrospective analyses that create this fragility while separating efficiency gains from changes in firms’ conduct. Using only the pre-merger data, we develop a regression-based approach that leverages exogenous changes in market structure to forecast prices after the merger. Finally, we implement structural merger analysis and show how estimates from all approaches can be synthesized for comprehensive evaluation. This methodological integration uncovers a fundamental tension: merger-induced efficiency gains were limited or, if significant, offset by increased coordination among remaining firms.
@@ 1 Introduction
Merger evaluation remains one of the most challenging problems in antitrust policy because of the difficulties separating efficiency gains from increased market power. This literature has developed two distinct methodologies: retrospective analyses that examine post-merger outcomes by applying event study methods like difference-in-differences (DiD) to pre- and post-merger data (e.g., Kim and Singal, 1993; Focarelli and Panetta, 2003; Hastings, 2004; Ashenfelter and Hosken, 2010; Ashenfelter et al., 2015; Carlton et al., 2019), and prospective analyses that predict merger effects using only pre-merger data (e.g., Nevo, 2000; Nevo and Whinston, 2010). A few important exceptions to this classification (Baker, 1999; Pesendorfer, 2003; Davis, 2005; Miller and Weinberg, 2017; Igami and Sugaya, 2022; Bruegge et al., 2024) underpin our work. For a more comprehensive overview of mergers, see Whinston (2006); Kwoka, Jr. (2014); Asker and Nocke (2021) and Kaplow (2024).
@@
We aim to fill this gap by developing a unified approach that bridges these methodologies, demonstrating its value through an analysis of three major airline mergers: Delta-Northwest (2008), United-Continental (2010), and American-US Airways (2013). Our empirical framework operates along two distinct dimensions: analyses using pre- and post-merger data versus only pre-merger data and reduced-form versus structural approaches. This approach creates four complementary analyses (see Table 1) that provide a comprehensive merger evaluation while revealing insights that any single approach would miss.
@@
Reduced-Form Retrospective Analysis Market Structure Analysis
@@
Our paper follows a systematic analysis that sequentially integrates reduced-form and structural approaches, assuming that mergers and market structure are exogenous. [^1] This assumption is nearly universal in the empirical merger literature, underlying both retrospective analyses and structural merger simulations, with notable exceptions being Dafny (2009); Aguirregabiria and Ho (2012); Ciliberto et al. (2021); Li et al. (2022), and Bontemps et al. (2023) among others. We begin with a retrospective analysis based on the DiD method using pre- and post-merger data (Section 3), then evaluate the robustness of this approach concerning the time trends to capture possibly changing demand. To explore these findings and make explicit the economic assumptions underlying the DiD framework, we then develop a structural framework of demand and supply such that the implied reduced-form price equation corresponds to the DiD estimation equation (Section 4). Our approach also clarifies assumptions implicit in DiD about airline pricing to determine merger effects.
Next, we consider the scenario typically faced by antitrust authorities and courts: evaluating mergers using only pre-merger data. In particular, we develop a regression-based approach that relies on the exogenous variation in market structure to forecast merger effects (Section 5). Finally, we implement a structural merger simulation that uses estimated effects from earlier sections to determine post-merger marginal costs and evaluate mergers under different assumptions of strategic conduct (Section 6).
The US airline industry provides an ideal setting for our analysis. It offers rich, comprehensive data spanning ticket prices, passenger numbers, seat capacity, and flight schedules from multiple authoritative sources: the Department of Transportation’s Origin and Destination Survey for ticket prices and passenger numbers, the Bureau of Transportation Statistics’ T100 database for seat capacity, and the Official Airline Guide (OAG) data for flight schedules. The industry’s extensive scrutiny in both academic research (e.g., Borenstein, 1990; Kim and Singal, 1993; Olley and Town, 2018; Carlton et al., 2019; Das, 2019; Li et al., 2022; Eizenberg and Zvuluni, 2024; Olsen et al., 2024; Remer and Orchinik, 2024) and legal proceedings (e.g., U.S. District Court, MA, 2023) provides crucial institutional context and validation points for our methodology.
@@
Next, in Section 4, we develop an equilibrium model with a linear market-level demand and supply relationship that links pricing behavior to mergers. In the process, we also derive the DiD estimation equation from Section 3 as a reduced-form expression of our structural pricing equation, explicitly delineating the economic assumptions that underpin the DiD strategy. To this end, we build on Bresnahan (1982) and propose a structural model that allows merger-specific coordinated effects (increased risk of coordinated behavior among rivals, see, for example, Porter (2020) and Loertscher and Marx (2021)), treating air travel as a homogeneous product with market-level quantity-weighted average prices.
@@
Our equilibrium model uncovers three critical insights in the implicit assumptions of standard retrospective analyses. First, consistent DiD estimation requires that mergers not alter firm conduct and that time-varying unobservable “demand shocks” evolve similarly across treated and control markets. Second, the error term in the DiD specification contains unobservable demand and supply shocks. While the cost shocks may reasonably be uncorrelated with merger decisions, demand shocks could systematically differ between treated and control markets—for instance, if firms are more likely to merge in markets with declining demand—and biasing the DiD estimates. Third, our estimates suggest that mergers affect firm conduct differently across market structures, indicating both unilateral (increase in market power of the merged firm) and coordinated effects that the retrospective analysis might miss.
@@
Next, in Section 5, we develop a regression-based approach that leverages exogenous variations in market structure to predict post-merger effects. Building on methodology pioneered by Baker (1999), this approach analyzes how prices vary with different market structures in the pre-merger period to forecast merger-induced changes. The variation in market structure offers insights into how similar structural changes might affect post-merger outcomes. We find that transitions from duopoly to monopoly are associated with modest but consistent fare increases (ranging from 2.9% to 6.2%) and substantial capacity reductions (15.7% to 27.6%). Conversely, markets transitioning from monopoly to duopoly show symmetric effects: fare decreases of 3.0% to 7.7%, and large capacity increases of 31.5% to 38.5%. While intuitive and simple, this model also captures efficiency induced by past changes in market structure. When a market evolved from a triopoly to a duopoly, the remaining airlines likely optimized their operations in response—and those efficiencies are present in the observed price changes.
@@
Finally, in Section 6, we implement an equilibrium-based merger simulation (Nevo, 2000) that considers different scenarios for post-merger cost efficiencies and conduct. In particular, we consider a random utility discrete-choice model of demand with differentiated products and Bertrand competition where firms choose prices. Following Verboven (1996), we employ a nested logit demand model to distinguish between nonstop and connecting services, allowing for rich substitution patterns across products and carriers. Our merger simulation analysis reveals that predicted price effects depend critically on assumptions about post-merger cost efficiencies and firms’ conduct. Specifically, we explore how post-merger prices depend on efficiency gains and changes in competitive conduct. Under standard competition assumptions and maximum efficiency gains (where merged entities adopt the lower costs of merging firms), the model predicts price decreases ranging from 2.9% to 5.1%.
Our bridging framework reveals a fundamental tension in airline merger evaluation that traditional approaches miss: given that the DiD estimates imply the absence of price effects, either efficiency gains were modest, or if significant, they were offset by increased coordination among remaining airlines. This finding creates a strategic dilemma for merging parties: claiming large efficiency gains requires acknowledging increased market power while arguing for limited competitive effects undermines the efficiency justification for the merger.
In summary, by integrating retrospective and prospective analysis, our analysis reveals how stable post-merger prices mask offsetting changes in efficiency and market conduct—a dynamic that isolated methodologies would miss. This integrated framework provides antitrust authorities with a more complete toolkit for evaluating the complex tradeoffs inherent in merger analysis.
@@
Our paper’s contributions to the literature are both methodological and empirical. Building on the extensive merger analysis literature cited throughout the paper, we integrate existing reduced-form retrospective and structural prospective analyses, drawing insights from each while addressing their limitations. Specifically, we demonstrate how retrospective studies that use post-merger data to measure price effects (Ashenfelter and Hosken, 2010) and prospective approaches that use pre-merger data and focus on structural estimation of efficiency gains (Farrell and Shapiro, 1990) can be integrated to provide more robust merger analysis, under the assumption of exogenous mergers.
@@ 2 Data and Descriptive Analysis
We focus on three major mergers in the airline industry. Delta and Northwest Airlines announced their merger plan on April 14, 2008; the shareholders approved it on September 26, 2008, and the US Department of Justice (DOJ) approved it on October 29, 2008. United and Continental announced their merger plan on May 2, 2010, gaining shareholder approval on September 17, 2010. The merger was completed on October 1, 2010, forming United Continental Holdings, although the full operational integration was completed only in mid-2012. The merger between American Airlines and US Airways was announced on February 14, 2013. This merger faced initial opposition from the DOJ, which filed a lawsuit on August 13, 2013, to block the proposed merger. However, a settlement was reached on November 12, 2013, allowing the merger to be completed on December 9, 2013.333While Kim and Singal (1993) examine price effects at both merger announcement and implementation (approval) dates, our analysis focuses exclusively on the pre- and post-approval periods. This approach allows us to focus on the real market effects following regulatory clearance rather than the anticipatory effects, providing a more focused examination of realized competitive outcomes.
@@
Next, we discuss the selection of “treated” and “control” markets, an approach the US court system accepts when evaluating merger cases. [^4] In our analysis, we also consider the sample of markets connecting the top 50 Metropolitan Statistical Areas (MSAs), a standard approach widely used in the literature. For example, Berry and Jia (2010) examine airports located in medium-to-large metropolitan areas with at least 850,000 people in 2006, which roughly corresponds to the top 50 MSAs. For each merger, we classify a city pair market as a “treated” market if both merging parties provided nonstop service before the merger. [^5] A carrier is considered ‘nonstop’ in a city route if it conducts 10 nonstop operations (five round-trips). Furthermore, we separately focus on 2-to-1 markets, the duopoly markets, where only the merging parties provide nonstop service, and 3-to-2 markets, the triopoly markets, where one other carrier provides nonstop service.
@@
Likewise, we classify a market as a “control” market if the merger does not affect its market structure. In other words, in such a market, only one carrier can be the merging party, so if it is a duopoly or a triopoly before the merger, it remains the same after the merger. Thus, the control markets have the same number of nonstop carriers as the overlap routes in the pre-merger period. As with the treated market, we also restrict our attention to control markets where, whenever applicable, the merging party does not have a substantial connecting presence.
@@ 3 Retrospective Analysis
We use the event-study approach to separately assess the impact of three legacy airline mergers (Delta-Northwest, United-Continental, and American-US Airways) on three key economic variables: average fares, passenger volume, and seat availability.
The analysis compares the mean of each variable in control markets before and after a merger and compares that difference with the difference in means in treated markets. In particular, we estimate the following regression specification to determine the average treatment on the treated markets under the DiD assumptions:
@@ 3.1 Estimation Results
Our parameter of interest is $\beta^{\texttt{DiD}}$, which, under the parallel trends assumptions, is the effect of a merger on the outcome variables in the treated markets. The estimation results from (1) are shown in Table 3. Panel A presents estimates using the (mean deviation) fixed-effects method. We estimate that the average prices decreased by 4.02% (DL-NW), 3.54% (UA-CO), and 7.50% (AA-US). The passenger volume increased by 2.84% (DL-NW), 5.87% (UA-CO), and 9.19% (AA-US). [^6] This estimate is a weighted average of the parameter estimates across markets and time (de Chaisemartin and D’Haultfœuille, 2020) and should be interpreted as a percentage change in the outcome variable. The percentage change for $\beta^{\texttt{DiD}}$ is given by $(\exp(\beta^{\texttt{DiD}})-1)\times 100$ (Halvorsen and Palmquist, 1980). Almost all of these estimates are statistically significant at the 1%-level, except that the estimated effect on passenger volume following the DL-NW merger is noisy. Most notably, the number of offered seats, i.e., the allocated capacities, expanded by 24.11% (DL-NW), 29.44% (UA-CO), and 16.53% (AA-US)–all statistically significant.
@@ 3.2 Robustness Analysis
The most natural and simple way to check the robustness is to run first differences instead of fixed-effects, as done in Kim and Singal (1993). The first difference and fixed-effect methods should yield similar estimates. [^8] They yield the same estimates if there is only one period before and one period after the mergers. However, they differ in their identifying assumptions, which in the presence of time trends may result in different estimates (Wooldridge, 2010). In particular, the fixed-effects method requires strict exogeneity, i.e., $\text{Treated}_{mt}\times\text{Post-Merger}_{mt}$ is uncorrelated with $\varepsilon_{mt}$ across all periods. In contrast, the first differences require that only the variable is uncorrelated with the change in $\varepsilon_{mt}$ immediately before and after the merger. The variable $\text{Treated}_{mt}\times\text{Post-Merger}_{mt},$ changes value only in a single quarter, suggesting that the first differences assumption is less demanding than the fixed-effect method. Furthermore, in a DiD framework, the first differences have an additional advantage: the constant term captures any differential linear trends between treatment and control markets, slightly relaxing the parallel trends assumption.
The results from using the first differences to estimate (1) are shown in Table 3 Panel B. The contrast with our fixed-effects estimates is striking in three key dimensions.
First, while fixed-effects estimation suggests that mergers lead to a 5.6% decrease in prices, the first-differences approach suggests prices increase slightly. However, they are measured with noise, except for the UA-CO merger, for which the price increases by 4.5%. Second, the effects of the merger on passenger volume are significantly different. The DL-NW merger shows a substantial 21.8% decrease in passengers using first differences, compared to a small positive effect using the fixed-effects method. Only AA-US maintains a positive passenger effect (27.5% increase) in first differences. Third, while capacity effects remain positive and significant across both specifications, their magnitude is notably smaller under the first differences approach. For example, the UA-CO capacity effect drops from 25.8% to 3.7%.
The first-differences method implicitly allows for different time trends between treatment and control groups through the constant term. In the airline context, such differential trends could reflect market-specific changes in local economic conditions or evolving travel patterns. The differences in the estimates using the two methods highlight the importance of capturing time trends for merger analysis.
The stark differences between fixed-effects and first-differences estimates raise concerns about whether the parallel trends assumption—fundamental to DiD analysis—holds in our setting. To investigate this issue more rigorously, we conduct pre-trend analyses using the event study approach.
@@
Our pre-trend analysis in Figure 2 indicates that the parallel trends assumption critical to DiD methodology is violated in this context. One approach to this violation is including market-specific time trends in our model. In airline markets, such trends reflect gradual changes in local economic conditions or evolving travel patterns, specifically on the demand side. However, incorporating these trends presents methodological challenges, particularly in choosing the appropriate functional form.
@@
Table 4 presents these results. Incorporating time trends alters the main findings in Table 3 Panel A: price decreases, and passenger volume increases either reverse or become statistically insignificant. This pattern holds across different trend specifications, suggesting the initial findings may have been driven by pre-existing market trajectories rather than merger effects.
@@
While our linear trend specification imposes strong assumptions about market evolution, specifically that trends are linear and homogeneous within groups, we validate our findings using more flexible methods. In Appendix B.6, we demonstrate that the synthetic DiD approach, which better captures the complex dynamics of airline markets, produces qualitatively similar results. This consistency across methodologies strengthens our confidence in the robustness of our findings. [^10] Specifically, the estimates from synthetic DiD are in Appendix B.6 Table B.5 Panel B and are similar to the first-difference results reported in Table 3 Panel B for prices and passenger volume. Our findings are similar to those of Remer and Orchinik (2024), who use the synthetic DiD method and find that mergers lead to higher prices.